# Quantifying non-Markovianity via local quantum Fisher information

### v2

@@ I Introduction
Over recent years, quantum mechanics has evolved from being a foundational theory in physics to a driving force behind numerous technological revolutions, bridging the gap between theoretical insights and experimental advancements. These quantum technologies, such as quantum computing and quantum communication, have garnered increasing interest due to their revolutionary potential [1, 2]. However, these technologies face significant challenges due to the inevitable interactions with the environment that lead to a loss of quantum properties through decoherence phenomena [3, 5, 4]. Weak coupling between the system and its environment characterises the evolution of these quantum systems [6, 7, 8]. In such weak couplings, it is often reasonable to assume that they are memoryless, i.e. no information flows back into the system later. such that the system dynamics is easier to handle and describe by the Markov and Born approximations hold, which allows us to obtain the time-local master equation within in the Lindblad form [9, 10], while in some cases the Markovian approximation breaks down, such as strong coupling between system and environment and non vanishing initial system environment correlation. In addition, it has been found that non-Markovianity can lead to a significant variety of physical effects in the dynamics of open quantum systems [15, 11, 12, 13, 14] and can serve as a resource in information theory. If the dynamics are of a Markovian nature, the loss of information will only occur in one direction flow of information from the system to the environment. Conversely, non-Markovian dynamics presents more interesting phenomena due to the memory effect and has been used in various quantum processes. Thus, the problem of the characterization of our non-Markovian effect quickly became a major topic in the theory of open quantum dynamics [16, 17, 18, 19, 73].
@@
The search for detecting non-Markovianity have produced an various measures, some of which we have already cited. These non-Markovian measures are equivalent for single-channel open two-level systems [45, 56]. However, they may not be exactly equivalent for multi-channel open two-level systems [57, 15]. A universal non-Markovian definition of open quantum systems remains elusive. In this study, we have presented a simple and mathematically viable approach based on the LQFI for measuring non-Markovianity in OQSs. The LQFI proposed by Kim et al [58], which is defined as the minimization of the quantum Fisher information associated with a local observable, which indicates the maximum information achievable within one of the subsystems. However, the LQFI has considerable potential as a tool for understanding the effects of quantum correlations beyond entanglement and for improving the precision of quantum estimation protocols [59]. In addition to the derivation of a new measure based-LQFI, the focus of our extension study is to compare it with the based on local quantum uncertainty measure for the detection of non-Markovianity. The so-called local quantum uncertainty is a measure of quantum correlations proposed by Girolami et al [60], which quantifies the minimum uncertainty resulting from applying local measurements to a subset of the quantum state using the concept of Wigner-Yanase information [61]. In addition to the importance of LQU as a quantitative measure of quantum correlations, He et al.[47] used it to derive a new measure to detect non-Markovianity. The LQU-based measure provides a clear physical interpretation and captures non-Markovianity in bipartite systems. The quantum uncertainty of local observables in the other free subsystem increases (decreases) in proportion to a non-Markovian (Markovian) process in the other subsystem. Recognizing that both LQU and LQFI share the crucial characteristic of measuring quantum correlations, as evidenced in numerous studies (e.g., Refs. [5, 62, 63]), we will leverage LQFI and its relationship to LQU within the context of quantum correlations to construct our measure for detecting non-Markovianity in open quantum systems. Our previous work [64] demonstrated the clear superiority of LQFI over LQU in any qubit-qudit system. A fundamental question arises: Can LQFI-based and LQU-based measures be compared in quantifying the degree of non-Markovianity, and which would be more effective in this context?
@@
In this paper, An alternative measure of non-Markovianity based on the LQFI is presented and contrasted with the measure LQU. Focusing on a bipartite system $AB$, where $A$ is the primary system interacting with the environment and $B$ is the ancilla, we assess the LQFI by applying local unit operations to the ancilla, effectively using it as a measuring device. In this context, the non-Markovianity of the system evolution is quantified by evaluating the non-monotonic tendencies of the LQFI. Furthermore, it is imperative to highlight that the measure of non-Markovianity derived by the LQFI shows a qualitative consistency with several other established measures used in the context of OQS. In our investigation, we have shown that this LQFI-based measure is consistent with metrics based on the divisibility concept, relative entropy and quantum mutual information and particular LQU. The structure of this manuscript unfolds as follows: Section II begins with a review of the concept of LQFI and local quantum uncertainty. Following this, a non-Markovianity measure based on LQFI for OQSs is introduced. Moving on to section III, we present illustrative examples to demonstrate the practical utility of this measure, juxtaposing it with the LQU-based measure. Finally, our findings are summarized in section IV.
@@ II.1 Non-Markovianity via local quantum uncertainty
LQU refers to the fundamental unpredictability of quantum phenomena within a local region of a quantum system. This reflects the intrinsically probabilistic nature of quantum behavior, and also highlights the limitations on the precision of simultaneous measurements of certain properties of quantum particles. This discord-type measure of quantum correlations $\mathcal{U}$, proposed by Girolami et al.[60], is defined as the minimum skew information [61] from the perspective of local measurements, in any bipartite quantum state. It is explicitly defined as
@@ II.2 Non-Markovianity via local quantum Fisher information
Local quantum Fisher information is a fundamental concept in quantum information theory and quantum estimation theory. Recently, it has emerged as an important quantifies the quantum correlations. It is defined as the minimum quantum Fisher information overall local Hamiltonian $L_{A}$ acting in the subspace of party $A$ of the bipartite system $AB$, i.e.,

### v3

@@ I Introduction
Over recent years, quantum mechanics has evolved from being a foundational theory in physics to a driving force behind numerous technological revolutions, bridging the gap between theoretical insights and experimental advancements. These quantum technologies, such as quantum computing and quantum communication, have garnered increasing interest due to their revolutionary potential [1, 2]. However, these technologies face significant challenges due to the inevitable interactions with the environment that lead to a loss of quantum properties through decoherence phenomena [3, 5, 4]. Weak coupling between the system and its environment characterises the evolution of these quantum systems [6, 7, 8]. In such weak couplings, it is often reasonable to assume that they are memoryless, i.e., no information flows back into the system later, such that the system dynamics is easier to handle and describe by the Markov and Born approximations hold, which allows us to obtain the time-local master equation within in the Lindblad form [9, 10], while in some cases the Markovian approximation breaks down, such as strong coupling between system and environment and non vanishing initial system environment correlation. In addition, it has been found that non-Markovianity can lead to a significant variety of physical effects in the dynamics of open quantum systems [15, 11, 12, 13, 14] and can serve as a resource in information theory. If the dynamics are of a Markovian nature, the loss of information will only occur in one direction flow of information from the system to the environment. Conversely, non-Markovian dynamics presents more interesting phenomena due to the memory effect and has been used in various quantum processes. Thus, the problem of the characterization of our non-Markovian effect quickly became a major topic in the theory of open quantum dynamics [16, 17, 18, 19, 65].
@@
The search for detecting non-Markovianity has produced various measures, some of which we have already cited. These non-Markovian measures are equivalent for single-channel open two-level systems [45, 56]. However, they may not be exactly equivalent for open two-level multichannel systems [57, 15]. A universal non-Markovian definition of open quantum systems remains elusive. In this study, we have presented a simple and mathematically viable approach based on the LQFI for measuring non-Markovianity in OQSs. The LQFI proposed by Kim et al [58], which is defined as the minimization of the quantum Fisher information associated with a local observable, indicates the maximum information achievable within one of the subsystems. However, the LQFI has considerable potential as a tool for understanding the effects of quantum correlations beyond entanglement and for improving the precision of quantum estimation protocols [59]. In addition to the derivation of a new LQFI-based measure, the focus of our extension study is to compare it with the local quantum uncertainty measure for the detection of non-Markovianity. The so-called local quantum uncertainty is a measure of quantum correlations proposed by Girolami et al [60], which quantifies the minimum uncertainty resulting from applying local measurements to a subset of the quantum state using the concept of Wigner-Yanase information [61]. In addition to the importance of LQU as a quantitative measure of quantum correlations, He et al.[47] used it to derive a new measure to detect non-Markovianity. The LQU-based measure provides a clear physical interpretation and captures non-Markovianity in bipartite systems. The quantum uncertainty of local observables in the other free subsystem increases (decreases) in proportion to a non-Markovian (Markovian) process in the other subsystem. Recognizing that both LQU and LQFI share the crucial characteristic of measuring quantum correlations, as evidenced in numerous studies (e.g., Refs. [5, 62, 63]), we will leverage LQFI and its relationship to LQU within the context of quantum correlations to construct our measure for detecting non-Markovianity in open quantum systems. Our previous work [64] demonstrated the clear superiority of LQFI over LQU in any qubit-qudit system. A fundamental question arises: Can LQFI-based and LQU-based measures be compared in quantifying the degree of non-Markovianity, and which would be more effective in this context?
@@
In this paper, An alternative measure of non-Markovianity based on the LQFI is presented and contrasted with the measure LQU. Focusing on a bipartite system $AB$, where the primary system $A$ is interacting with the environment and $B$ is the ancilla, we assess the LQFI by applying local unit operations to the ancilla, effectively using it as a measuring device. In this context, the non-Markovianity of the system evolution is quantified by evaluating the non-monotonic tendencies of the LQFI. Importantly, our measure operates within the framework of completely positive and trace-preserving (CPTP) dynamics, as it relies on the inclusion of an ancillary system, which inherently ties it to the CP-divisibility paradigm (see Fig.4 of Ref.[65] for details). This contrasts with measures based solely on positive (P)-divisibility, such as those based on trace distance [16], which are necessary and sufficient to detect non-Markovianity in P-indivisible channels but do not require an ancillary extension. Although P- and CP-divisibility represent distinct approaches to characterizing non-Markovianity, with P-divisibility capturing information backflow in isolated systems and CP-divisibility accounting for correlations in extended systems, our study focuses exclusively on the latter. By focusing on CPTP dynamics, we ensure consistency with ancilla-assisted metrics and maintain a physically meaningful description of open quantum system evolution. Furthermore, it is imperative to highlight that the measure of non-Markovianity derived by the LQFI shows a qualitative consistency with several other established measures used in the context of OQS. In our investigation, we have shown that this LQFI-based measure is consistent with metrics based on the divisibility concept, relative entropy and quantum mutual information and particular LQU. Notably, this alignment holds for CPTP dynamics, as our ancilla-dependent method naturally aligns with CP-divisibility conditions.
The structure of this manuscript unfolds as follows: Section II begins with a review of the concept of LQFI and local quantum uncertainty. Following this, a non-Markovianity measure based on LQFI for OQSs is introduced. Moving on to section III, we present illustrative examples to demonstrate the practical utility of this measure, juxtaposing it with the LQU-based measure. Finally, our findings are summarized in section IV.
@@ II.1 Non-Markovianity via local quantum uncertainty
Local quantum uncertainty (LQU) refers to the fundamental unpredictability of quantum phenomena within a local region of a quantum system. This reflects the intrinsically probabilistic nature of quantum behavior, and also highlights the limitations on the precision of simultaneous measurements of certain properties of quantum particles. This discord-type measure of quantum correlations $\mathcal{U}$, proposed by Girolami et al.[60], is defined as the minimum skew information [61] from the perspective of local measurements, in any bipartite quantum state. It is explicitly defined as
@@ II.2 Non-Markovianity via local quantum Fisher information
Local quantum Fisher information is a fundamental concept in quantum information theory and quantum estimation theory. Recently, it has emerged as an important quantifier of quantum correlations. It is defined as the minimum quantum Fisher information overall local Hamiltonian $L_{A}$ acting in the subspace of party $A$ of the bipartite system $AB$, i.e.,
@@ IV Concluding Remarks and Outlook
Building on these results, it is interesting to investigate and explain the experimental feasibility of detecting the CPTP non-Markovianity using LQFI. Its link with measurable quantities in quantum metrology and its precise relationship with quantum correlations are advantageous. This will enable us to recognize the limitations of optimization-based measurements and explore the avenues described above, including the potential role of LQFI and the need for alternative, more experimentally suitable approaches. Addressing this open question is crucial to bridge the gap between theoretical understanding and practical applications of non-Markovian quantum information processing.